OpenStudy (anonymous):

What is the slope of the line y = -3x + 6?

jimthompson5910 (jim_thompson5910):

hint: y = mx+b has a slope of m

jimthompson5910 (jim_thompson5910):

Example y = 2x + 3 is in the form y = mx+b where m = 2 so the slope of y = 2x+3 is m = 2